1.For a body dropped from a height, h, the speed, V,with which it strikes the ground depends only on h and g and a dimensionless numerical factor. Obtain an expression connecting v, h and g.



1
Expert's answer
2021-01-07T12:15:32-0500

"h=\frac{gt^2}{2},"

"v=gt, ~t^2=\frac{v^2}{g^2},"

"h=\frac{gv^2}{2g^2}=\frac{v^2}{2g},"

"v=\sqrt{2gh}."